# Addressing a fundamental limitation in deep vision models: lack of spatial attention

### v3

@@ 2 Related work
The closest concept to this work is event cameras [1], which aim to process only the content that has changed at the hardware level by using specially designed cameras. In video surveillance, some systems efficiently detect and focus on frames with notable movement or changes, enhancing surveillance effectiveness. However, while these approaches skip frames, they still process the entire image when they do choose to process a frame. In contrast, the proposed approach processes only specific regions of the image, providing finer granularity compared to previous methods, although it requires some memory to save previous outputs.
@@ 4 Potential solution II
The second method involves semantic segmentation, often regarded as a supertask. This is because the capability to label each individual pixel can generalize to other tasks, such as object recognition and object detection. (Instance segmentation, however, is even more comprehensive since it distinguishes between instances of the same object.) The idea here is to feed only the modified regions into the segmentation model. A benefit of some models, especially those based on U-Net architectures [9], is that they can process inputs of any resolution. After obtaining the output, it can be accurately positioned in the prior output map.
@@ 6 Appendix
Convolution and pooling modules of CNNs are shown in Figures 89, and  10.

### v4

@@ 2 Related work
The closest concept to this work is event cameras [1], which aim to process only the content that has changed at the hardware level by using specially designed cameras (See also [9]). In video surveillance, some systems efficiently detect and focus on frames with notable movement or changes, enhancing surveillance effectiveness. However, while these approaches skip frames, they still process the entire image when they do choose to process a frame. In contrast, the proposed approach processes only specific regions of the image, providing finer granularity compared to previous methods, although it requires some memory to save previous outputs.
@@ 4 Potential solution II
The second method involves semantic segmentation, often regarded as a supertask. This is because the capability to label each individual pixel can generalize to other tasks, such as object recognition and object detection [^6] Instance segmentation, however, is even more comprehensive since it distinguishes between instances of the same object.. The idea here is to feed only the modified regions into the segmentation model. A benefit of some models, especially those based on U-Net architectures [10], is that they can process inputs of any resolution. After obtaining the output, it can be accurately positioned in the prior output map. Please see Fig. 7.
@@
Figure 7: The schematic diagram of the second proposed solution highlights that only the modified regions (bounded by boxes fitting the changed areas, such as the top-left region or altered grid locations) are processed by U-Net. The segmentation results for these regions are then merged into the previous output map.
@@ 6 Appendix
Convolution and pooling modules of CNNs are shown in Figures 910, and  11.